OpenStudy (anonymous):

Write an explicit formula for the sequence determined by this recursion formula: t{1} = 1/2; t{n} = t{n-1} + 1/n(n+1)

OpenStudy (anonymous):

t{1} = 1/2; t{n} = t{n-1} + 1/n(n+1)

OpenStudy (amistre64):

t{1} = 1/2 = 1/2 = 1/2 t{2} = 1/2 + 1/2(3) = 4/6 = 2/3 t[3] = 4/6 + 1/3(4) = 9/12 = 3/4 t{4} = 9/12 + 1/4(5) = 48/60 = 4/5 i see a pattern

OpenStudy (amistre64):

\[t_{n} = \frac{n}{n+1}\]

OpenStudy (anonymous):

ohhh, so you first find the pattern?

OpenStudy (amistre64):

pattern finding is not the ideal way to do it; but it is a simple way alot of the times

OpenStudy (amistre64):

there is a more analytical method which is long and drawn out

OpenStudy (anonymous):

ohh, alright.

OpenStudy (anonymous):

what is it?

OpenStudy (amistre64):

it involves working the recurrsion back to a{1} and applying summations along the way to catch all the details

OpenStudy (amistre64):

well, in this case t{1}

OpenStudy (amistre64):

\[ \begin{align}t{n} &= t_{n-1} + \frac{1}{n(n+1)}\\\\ t_{n-1}& = t_{n-2} + \frac{1}{(n-1)((n-1)+1)}\\\\ &\text{substitute in for }t_{n-1}\\\\ t_{n}& = \left(t_{n-2} + \frac{1}{(n-1)((n-1)+1)}\right)+ \frac{1}{n(n+1)}\\\\ & = t_{n-2} + \frac{1}{(n-1)((n-1)+1)}+ \frac{1}{n(n+1)}\\\\ \end{align}\] you continue in this substituing for lower and lower t{n-r} till you can restructure it

OpenStudy (anonymous):

I don't understand what you just inputted.. o.o

OpenStudy (amistre64):

\[t_{n-2}=t_{n-3}+\frac{1}{(n-1)((n-2)+1)}\] thats because you are not familiar with the process as written, even tho you are familiar with replacing like values

OpenStudy (amistre64):

im just rewriting the original recurrsion equation in terms of lesser and lesser terms

OpenStudy (amistre64):

t{n-2} comes before t{n-1} which comes before t{n} .... just re writting the original in an order of terms that come before it

OpenStudy (amistre64):

\begin{align}t{n} &= t_{n-1} + \frac{1}{n(n+1)}\\\\ t_{n-1}& = t_{n-2} + \frac{1}{(n-1)((n-1)+1)}\\\\ &\text{substitute in for }t_{n-1}\\\\ t_{n}& = \left(t_{n-2} + \frac{1}{(n-1)((n-1)+1)}\right)+ \frac{1}{n(n+1)}\\\\ & = t_{n-2} + \frac{1}{(n-1)((n-1)+1)}+ \frac{1}{n(n+1)}\\\\ t_{n-2}&=t_{n-3}+\frac{1}{(n-2)((n-2)+1)}\\\\ t_{n}& = \left(t_{n-3}+\frac{1}{(n-2)((n-1)} \right)+ \frac{1}{(n-1)((n)}+ \frac{1}{n(n+1)}\\\\ & = t_{n-3}+\frac{1}{(n-2)(n-1)}+ \frac{1}{(n-1)n}+ \frac{1}{n(n+1)}\\\\ & = t_{n-r}+\frac{1}{(n-(r-1))(n-(r-2))}+ \frac{1}{(n-(r-2))(n-(r-1))}+ ...\\\\ &\text{when r=n-1; t{n-r}=t{1}}\\\\ & = t_{1}+\frac{1}{(2)(3)}+ \frac{1}{(3)(2)}+ \frac{1}{(4)(3)} +...+ \frac{1}{(n)(n+1)}\\\\ \end{align} the trick then becomes to find a summation of that tail end since we have defined it in terms of t{1}

OpenStudy (amistre64):

but like i said, long and drawn out process

OpenStudy (amistre64):

if you can find a pattern that works, that tends to be the simplest way to go